OpenStudy (haltiamali):

Simplify (-3)^2(x)(-x) (the two is an exponent) Please explain? Final tomorrow!

OpenStudy (haltiamali):

\[-3^{2}(x)(-x)\]

OpenStudy (anonymous):

\[(-3)^2=3^2\]

OpenStudy (anonymous):

(-3)^2 = 9 actually

OpenStudy (anonymous):

-9x